The diameter of Earth is about twice that of Mars. The diameter of Mars is about three times that of Pluto. Together, the sum of the diameters of Earth, Mars, and Pluto is about 30,000 kilometers less than the diameter of Uranus. If the diameter of Uranus is about 52,000 kilometers, what equation would be used to find the approximate diameter of Pluto? Let d = diameter of Pluto.
Answer by ankor@dixie-net.com(12692)

The diameter of Earth is about twice that of Mars.
E = 2M
:
The diameter of Mars is about three times that of Pluto.
M = 3d & E = 2M
Therefore:
E = 2(3d)
E = 6d
:
Together, the sum of the diameters of Earth, Mars, and Pluto is about 30,000 kilometers less than the diameter of Uranus.
E + M + d = U - 30000
:
If the diameter of Uranus is about 52,000 kilometers, what equation would be used to find the approximate diameter of Pluto? Let d = diameter of Pluto.
d = U - 30000 - E - M
d = 52000 - 30000 - E -M
d = 22000 - E - M
:
Substitute 6d for E and 3d for M
d = 22000 - 6d - 3d
d = 22000 - 9d
d + 9d = 22000
10d = 22000
d = 22000/10
d = 2200